The height h (in feet) of an object that is dropped from the height of s feet is given by the formula h = s-16t^2, where t is the time the object has been falling. A 4 foot tall woman on a sidewalk looks directly overhead and sees a window washer drop a bottle from 6 story. How long does she have to get out of the way? Round to the nearest tenth. (A story is 12 feet.)

Answer by stanbon(75887)

The height h (in feet) of an object that is dropped from the height of s feet is given by the formula h = s-16t^2, where t is the time the object has been falling. A 4 foot tall woman on a sidewalk looks directly overhead and sees a window washer drop a bottle from 6 story. How long does she have to get out of the way? Round to the nearest tenth. (A story is 12 feet.)
--------------------
The object will hit the top of her head when it's height is 4 ft.
---------------------
EQUATION:
4 = 6*12-16t^2
4 = 72 -16t^2
16t^2 = 68
t^2 = 68/16
t = 2sqrt(17)/4 = sqrt(17)/2 seconds
t is approximately 2.06 seconds
=============
Cheers,
Stan H.